# Quantum entanglement in phase space

### v2

@@ Methods
Acknowledgments.– We are indebted to M. Walschaers for inspiring and insightful discussions on the topic and to L. H. Zaw and E. Agudelo for valuable comments on our work. This work is supported by the National Natural Science Foundation of China (Grants No. 12125402, No. 12405005, No. 12447157), the Beijing Natural Science Foundation (Grant No. Z240007), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), and the China Postdoctoral Science Foundation (No. 2023M740119 and 2024M760072). J. Guo acknowledges the CPSF Postdoctoral Fellowship Program (No. GZB20240027). MF was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by the Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
@@
I I. Criterion (I) can lower bound the entanglement negativity
@@
$=\iint_{\infty}^{\infty}\text{d}x\text{d}pW_{S(\rho^{T}_{B})}(x,p,0,0)$
@@
In fact, the bound is tight for Schmidt-correlated states $\sigma=\sum_{nm}c_{nm}|n,n\rangle\langle m,m|$, where $c_{nm}$ are real coefficients. We will prove it in the following.
The partial trace on mode B yields $\sigma^{T_{B}}=\sum_{nm}c_{nm}|n,m\rangle\langle m,n|$. The norm of $\sigma^{T_{B}}$ is $\sigma^{T_{B}}||\equiv\sqrt{\left(\sigma^{T_{B}}\right)^{\dagger}\left(\sigma^{T_{B}}\right)}=\sqrt{\sum_{n,m}c^{2}_{nm}|m,n\rangle\langle m,n|}=\sum_{n,m}c_{nm}|m,n\rangle\langle m,n|$. By multiplying a swap operator $S_{\text{swap}}=\sum_{n,m}|m,n\rangle\langle n,m|$, we have $\sigma^{T_{B}}S_{\text{swap}}=\sum_{n^{\prime},m^{\prime},n,m}c_{nm}(|n,m\rangle\langle m,n|)$ $(|m^{\prime},n^{\prime}\rangle\langle n^{\prime},m^{\prime}|)=\sum_{n,m}c_{nm}|m,n\rangle\langle m,n|$. Thus it is found that $\sigma^{T_{B}}||=\sigma^{T_{B}}S_{\text{swap}}$.
@@
Here, we use the Wigner function of a swap operator $4\pi\delta\left(x_{1}-x_{2}\right)\delta\left(p_{1}-p_{2}\right)$. The last raw is the left-hand side of Criterion (I) under the transformation with a time-reversal $(x,p)\rightarrow(x,-p)$ and $\theta=\pi/4$.
II II. Choice of $\Delta=\pm 1$ in linear transformations
II.1 A. Criterion (I) requires $\Delta=-1$
@@
In the case of $\Delta=+1$, the quadratures ($x^{\prime},p^{\prime}$) is obtained from ($x,p$) through a unitary transformation, i.e. $x^{\prime}=U^{\dagger}xU,p^{\prime}=U^{\dagger}pU$. Then according to Eq. (37) and Eq. (38) of Sec. IV in SM, Criterion (I) can be expressed as
@@
where $h_{1},h_{2}$ are state transformations in Eqs. (35,36) and $U,U^{\prime}$ are unitary transformation on mode B and modes AB, respectively. After the unitary transformation, $\rho^{\prime}_{AB}$ should be a physical state, so that $B^{\prime}$ is also physical. Therefore, the Wigner function of the reduced mode $B^{\prime}:(X,P)$ is always upper-bounded by $1/(2\pi)$, meaning Criterion (I) will not be violated.
II.2 B. Criterion (III) requires $\Delta=+1$
@@
$\Delta=-1$ indicates a mirror reflection, as an example we consider the time reversal $(x,p)\rightarrow(x,-p)$. The local linear coordinate transformation corresponds to partial transpose of the state $W_{\rho}(x_{A},p_{A},x_{B},-p_{B})=W_{\rho^{T_{B}}}(x_{A},p_{A},x_{B},p_{B})$. Based on Eq. (44), we obtain
$W_{\operatorname{tr}_{-}(\rho^{T_{B}})}\left(x_{+},p_{+}\right)=$ $\iint_{-\infty}^{\infty}W_{\rho^{T_{B}}}\left(\frac{x_{+}+x_{-}}{\sqrt{2}},\frac{p_{+}+p_{-}}{\sqrt{2}},\frac{x_{+}-x_{-}}{\sqrt{2}},\frac{p_{+}-p_{-}}{\sqrt{2}}\right)\text{d}x_{-}\text{d}p_{-}$
$=$ $2\iint_{-\infty}^{\infty}W_{\rho^{T_{B}}}\left(x_{-},p_{-},-x_{-}+\sqrt{2}x_{+},-p_{-}+\sqrt{2}p_{+}\right)\text{d}x_{-}\text{d}p_{-}$
$=$ $2\iint_{-\infty}^{\infty}W_{\rho^{T_{B}}}\left(x,p,-x+x_{0},-p+p_{0}\right)\text{d}x\text{d}p$
$=$ $2\iint_{-\infty}^{\infty}W_{\rho}\left(x,p,-x+x_{0},p-p_{0}\right)\text{d}x\text{d}p$
$=$ $2\iint_{-\infty}^{\infty}W_{\mathbbm{1}\otimes f(\rho)}\left(x,p,x,-p\right)\text{d}x\text{d}p,$ (24)
where $f:(x,-p)\rightarrow(-x+x_{0},p-p_{0})$ corresponds to a local unitary.
In [12], the author pointed out that the negativity of the reduced Wigner function implies partial transpose negativity. For $\rho^{T_{B}}$, Theorem 2 in [12] gives
$W_{\operatorname{tr}_{-}(\rho^{T_{B}})}(x_{+},p_{+})\ngeq 0\Rightarrow\left(\rho^{T_{B}}\right)^{T_{B}}=\rho\nsucceq 0,$ (25)
which obviously conflicts with the fact that for any physical state $\rho\succeq 0$. Therefore, we obtain the following conclusion that
$W_{\operatorname{tr}_{-}(\rho^{T_{B}})}(x_{+},p_{+})\geq 0,$ (26)
which means Criterion (III) holds for any states.
III III. An intuitive phase-space counterpart to the Duan-Simon criterion and their comparison
@@
Observation S1. Consider first performing a transposition on mode B, followed by applying an arbitrary beam-splitter operation $S_{\mathrm{BS}}=\left(\begin{array}[]{cc}\cos(\theta)&\sin(\theta)\\ \sin(\theta)&-\cos(\theta)\end{array}\right)$ to the two modes, with the splitting ratio determined by the angle $\theta$. The coordinate transformations are given by $\left(x_{A}^{\prime},x_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(x_{A},x_{B}\right)^{T}$, $\left(p_{A}^{\prime},p_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(p_{A},-p_{B}\right)^{T}$. For any separable bipartite state $\rho_{AB}$ we have
@@
where $W^{\prime}\left(x_{B}^{\prime},p_{B}^{\prime}\right)=\iint_{-\infty}^{\infty}W_{AB}\left(x_{A},p_{A},x_{B},p_{B}\right)\text{d}x_{A}^{\prime}\text{d}p_{A}^{\prime}$ represents the state of one output mode after the beam-splitter transformation. Note that a transposition on mode B is not a physical operation, so that the measurements required are experimentally challenging and effectively equivalent to a full state tomography.
@@
By performing a partial transposition on mode $B$ and then sending the two modes through a beam-splitter $S_{\mathrm{BS}}=\left(\begin{array}[]{cc}\cos(\theta)&\sin(\theta)\\ \sin(\theta)&-\cos(\theta)\end{array}\right)$, which corresponds to the coordinate transformations $\left(x_{A}^{\prime},x_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(x_{A},x_{B}\right)^{T}$, $\left(p_{A}^{\prime},p_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(p_{A},-p_{B}\right)^{T}$, we obtain the Wigner function of one output mode of the beam-splitter,
$W_{B}^{\prime}(x_{B}^{\prime},p_{B}^{\prime})=\frac{1}{2\pi\left(-2c\sin(\theta)\cos(\theta)+m\cos^{2}(\theta)+n\sin^{2}(\theta)\right)}\exp\left(-\frac{x_{B}^{\prime 2}+p_{B}^{\prime 2}}{2\left(-2c\sin(\theta)\cos(\theta)+m\cos^{2}(\theta)+n\sin^{2}(\theta)\right)}\right).$ (32)
@@
IV IV. Detailed proof of criterion (I)
@@
Let $0\leqslant\theta<\pi$ be a constant angle, and let $(X,P)$ represent a constant coordinate. Then, define the following two coordinate transformations according to $(x^{\prime},p^{\prime})$, along with the corresponding state transformations $h_{1},h_{2}$,
@@
where in the last row we have applied the substitution $\left(x\rightarrow x+\frac{\sin\theta}{\cos\theta}X,p\rightarrow p+\frac{\sin\theta}{\cos\theta}P\right)$. Now, assume that $\rho_{AB}$ is a separable pure state, meaning $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$ should also correspond a physical state. Next, perform a beam-splitter transformation given by $\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{array}\right)$ on the two modes of $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$. The resulting Wigner function is given by
@@
V V. Criterion (I) is sufficient and necessary for Gaussian states
@@
$W_{AB}(\cos\theta x,\cos\theta p,\sin\theta x^{\prime},\sin\theta p^{\prime})$ (40)
$=$ $\frac{\exp\left(\frac{mt^{2}\cos^{2}(\theta)\left(-c_{1}^{2}p^{2}-c_{2}^{2}x^{2}+mn(p^{2}+x^{2})\right)+\sin(\theta)\left(n\sin(\theta)\left(-c_{1}^{2}p^{2}-c_{2}^{2}t^{4}x^{2}+mn(p^{2}+t^{4}x^{2})\right)+2c_{2}t\cos(\theta)\left(c_{1}c_{2}t^{2}x^{2}+mnp^{2}\right)\right)-c_{1}t\sin(2\theta)\left(c_{1}c_{2}p^{2}+mnt^{2}x^{2}\right)}{2t^{2}\left(c_{1}^{2}-mn\right)\left(mn-c_{2}^{2}\right)}\right)}{4\pi^{2}\sqrt{\left(c_{1}^{2}-mn\right)\left(c_{2}^{2}-mn\right)}}$
@@
$I=\int_{-\infty}^{\infty}W_{AB}(\cos\theta x,\cos\theta p,\sin\theta x^{\prime},\sin\theta p^{\prime})\text{d}x\text{d}p=\frac{1}{2\pi\sqrt{\frac{\left(m+nt^{2}+\cos(2\theta)(m-nt^{2})-2c_{1}t\sin(2\theta)\right)\left(mt^{2}+n+\cos(2\theta)(mt^{2}-n)+2c_{2}t\sin(2\theta)\right)}{4t^{2}}}}.$ (41)
@@
$t^{2}$ $=\frac{\left|c_{2}\right|m+\left|c_{1}\right|n}{\left|c_{1}\right|m+\left|c_{2}\right|n}$ (43)
$\tan(2\theta)$ $=\frac{-2\sqrt{\left|c_{2}\right|m+\left|c_{1}\right|n}\sqrt{\left|c_{1}\right|m+\left|c_{2}\right|n}}{(m-n)(m+n)}$
By simplifying the entanglement condition $I>1/(2\pi)$, we arrive precisely at Eq.(39).
VI VI. A comparison between criterion (III) and previous criteria based on Wigner functions
Reference [11] introduces an entanglement criterion based on reduced states. The Wigner function is first expressed in the $\left\{\hat{a}_{+},\hat{a}_{-}\right\}$ basis as $W_{\rho}\left(\alpha_{+},\alpha_{-}\right)$. By tracing out the $a_{+}$ mode, the reduced Wigner function is obtained as $W_{\text{tr}_{-}(\rho)}\left(\alpha_{+}\right)=\int\text{d}^{2}\alpha_{-}W_{\rho}\left(\alpha_{+},\alpha_{-}\right)$. If any point in the reduced Wigner function $W_{\text{tr}_{-}(\rho)}$ is negative, the state must be entangled. This criterion is a special case of our lower bound when choosing $(x^{\prime}=x_{0}-x,p^{\prime}=p_{0}-p)$, where $x_{0}=\sqrt{2}x_{+}$ and $p_{0}=\sqrt{2}p_{+}$ are constants. This is because

### v3

@@ Results
Acknowledgments.– We are indebted to M. Walschaers for inspiring and insightful discussions on the topic and to L. H. Zaw and E. Agudelo for valuable comments on our work. This work is supported by the National Natural Science Foundation of China (Grants No. 12125402, No. 12534016, No. 12405005, No. 12505010, No. 12447157), the Beijing Natural Science Foundation (Grant No. Z240007), and the China Postdoctoral Science Foundation (No. 2023M740119 and 2024M760072). J. Guo acknowledges the CPSF Postdoctoral Fellowship Program (No. GZB20240027). MF was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by the Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
@@
I. Criterion (I) can lower bound the entanglement negativity
@@
$=\iint_{-\infty}^{\infty}\text{d}x\text{d}pW_{S(\rho^{T}_{B})}(x,p,0,0)$
@@
In fact, the bound is tight for Schmidt-correlated states $\sigma=\sum_{nm}c_{nm}|n,n\rangle\langle m,m|$, where $c_{nm}$ are nonnegative real coefficients. We will prove it in the following.
The partial transpose on mode B yields $\sigma^{T_{B}}=\sum_{nm}c_{nm}|n,m\rangle\langle m,n|$. The norm of $\sigma^{T_{B}}$ is $\sigma^{T_{B}}||\equiv\sqrt{\left(\sigma^{T_{B}}\right)^{\dagger}\left(\sigma^{T_{B}}\right)}=\sqrt{\sum_{n,m}c^{2}_{nm}|m,n\rangle\langle m,n|}=\sum_{n,m}c_{nm}|m,n\rangle\langle m,n|$. By multiplying a swap operator $S_{\text{swap}}=\sum_{n,m}|m,n\rangle\langle n,m|$, we have $\sigma^{T_{B}}S_{\text{swap}}=\sum_{n^{\prime},m^{\prime},n,m}c_{nm}(|n,m\rangle\langle m,n|)$ $(|m^{\prime},n^{\prime}\rangle\langle n^{\prime},m^{\prime}|)=\sum_{n,m}c_{nm}|m,n\rangle\langle m,n|$. Thus it is found that $\sigma^{T_{B}}||=\sigma^{T_{B}}S_{\text{swap}}$.
@@
Here, we use the Wigner function of a swap operator $4\pi\delta\left(x_{1}-x_{2}\right)\delta\left(p_{1}-p_{2}\right)$. The last row is the left-hand side of Criterion (I) under the transformation with a time-reversal $(x,p)\rightarrow(x,-p)$ and $\theta=\pi/4$.
II. Choice of $\Delta=\pm 1$ in linear transformations
A. Criterion (I) requires $\Delta=-1$
@@
In the case of $\Delta=+1$, the quadratures ($x^{\prime},p^{\prime}$) are obtained from ($x,p$) through a unitary transformation, i.e. $x^{\prime}=U^{\dagger}xU,p^{\prime}=U^{\dagger}pU$. Then according to Eq. (40) and Eq. (41) of Sec. IV in SM, Criterion (I) can be expressed as
@@
where $h_{1},h_{2}$ are state transformations in Eqs. (38,39) and $U,U^{\prime}$ are unitary transformation on mode B and modes AB, respectively. After the unitary transformation, $\rho^{\prime}_{AB}$ should be a physical state, so that $B^{\prime}$ is also physical. Therefore, the Wigner function of the reduced mode $B^{\prime}:(X,P)$ is always upper-bounded by $1/(2\pi)$, meaning Criterion (I) will not be violated.
B. Criterion (III) requires $\Delta=+1$
@@
$\Delta=-1$ indicates a mirror reflection, as an example we consider the time reversal $(x,p)\rightarrow(x,-p)$. The local linear coordinate transformation corresponds to partial transpose of the state $W_{\rho}(x_{A},p_{A},x_{B},-p_{B})=W_{\rho^{T_{B}}}(x_{A},p_{A},x_{B},p_{B})$. Define
$I_{0}:=\iint_{-\infty}^{\infty}W_{\rho}(x,p,\,x,-p)\,\mathrm{d}x\,\mathrm{d}p.$ (23)
Substituting
$W(\bm{\xi})=\int\frac{\mathrm{d}^{4}\bm{\alpha}}{(2\pi)^{4}}\,\exp\left(-i\,\bm{\xi}^{\mathsf{T}}\Omega\,\bm{\alpha}\right)\chi(\bm{\alpha}),$ (24)
we obtain
$I_{0}=\int\frac{\mathrm{d}^{4}\bm{\alpha}}{(2\pi)^{4}}\,\chi(\bm{\alpha})\iint\mathrm{d}x\,\mathrm{d}p\;\exp\Big(-i\big[x(\alpha_{p_{A}}+\alpha_{p_{B}})-p(\alpha_{x_{A}}-\alpha_{x_{B}})\big]\Big).$ (25)
The integrals over $x$ and $p$ yield Dirac delta functions, $\int\mathrm{d}x\,e^{-ix(\alpha_{p_{A}}+\alpha_{p_{B}})}=2\pi\,\delta(\alpha_{p_{A}}+\alpha_{p_{B}})$,
$\int\mathrm{d}p\,e^{ip(\alpha_{x_{A}}-\alpha_{x_{B}})}=2\pi\,\delta(\alpha_{x_{A}}-\alpha_{x_{B}})$, so that
$I_{0}=\frac{1}{(2\pi)^{2}}\iint\mathrm{d}\alpha_{x}\,\mathrm{d}\alpha_{p}\;\chi(\alpha_{x},\alpha_{p},\,\alpha_{x},-\alpha_{p}).$ (26)
Using $\chi(\bm{\alpha})=\operatorname{Tr}\left[\rho\,e^{i\hat{\bm{\xi}}^{\mathsf{T}}\Omega\,\bm{\alpha}}\right]$, we can obtain $I_{0}=\operatorname{Tr}[\rho\,K]$, where
$K:=\frac{1}{(2\pi)^{2}}\iint\mathrm{d}\alpha_{x}\,\mathrm{d}\alpha_{p}\;\exp\Big(i\big[\alpha_{p}(\hat{x}_{A}-\hat{x}_{B})-\alpha_{x}(\hat{p}_{A}+\hat{p}_{B})\big]\Big).$ (27)
Define $\hat{X}_{-}:=\hat{x}_{A}-\hat{x}_{B}$ and $\hat{P}_{+}:=\hat{p}_{A}+\hat{p}_{B}$. Since $\hat{X}_{-}$ and $\hat{P}_{+}$ commute, the exponential factorizes as $e^{i(\alpha_{p}\hat{X}_{-}-\alpha_{x}\hat{P}_{+})}=e^{i\alpha_{p}\hat{X}_{-}}\,e^{-i\alpha_{x}\hat{P}_{+}}$, and
$K=\delta(\hat{X}_{-})\,\delta(\hat{P}_{+})=\delta(\hat{x}_{A}-\hat{x}_{B})\,\delta(\hat{p}_{A}+\hat{p}_{B}).$ (28)
Since $K\geq 0$, it follows that
$I_{0}=\operatorname{Tr}[\rho\,K]\geq 0.$ (29)
For a general $\Delta=-1$ transformation $M=SR$ where $S$ is symplectic and $R$ is the standard reflection, the corresponding operator is $K_{M}$ is unitarily equivalent to $K$. An additional displacement can be considered in a similar way. So the conclusion extends to all $\Delta=-1$ cases. Therefore, Criterion (III) holds for all states when $\Delta=-1$ and cannot be used to detect entanglement.
III. An intuitive phase-space counterpart to the Duan-Simon criterion and their comparison
@@
Observation S1. Consider first performing a transposition on mode B, followed by applying an arbitrary beam-splitter operation $S_{\mathrm{BS}}=\left(\begin{smallmatrix}\cos(\theta)&\sin(\theta)\\ \sin(\theta)&-\cos(\theta)\end{smallmatrix}\right)$ to the two modes, with the splitting ratio determined by the angle $\theta$. The coordinate transformations are given by $\left(x_{A}^{\prime},x_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(x_{A},x_{B}\right)^{T}$, $\left(p_{A}^{\prime},p_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(p_{A},-p_{B}\right)^{T}$. For any separable bipartite state $\rho_{AB}$ we have
@@
where $W_{B^{\prime}}\left(x_{B}^{\prime},p_{B}^{\prime}\right)=\iint\mathrm{d}x_{A}^{\prime}\mathrm{d}p_{A}^{\prime}W_{AB}^{\left(\mathrm{BS}\circ T_{B}\right)}\left(x_{A}^{\prime},p_{A}^{\prime},x_{B}^{\prime},p_{B}^{\prime}\right)$ represents the state of one output mode after the beam-splitter transformation. Note that a transposition on mode B is not a physical operation, so that the measurements required are experimentally challenging and effectively equivalent to a full state tomography.
@@
By performing a partial transposition on mode $B$ and then sending the two modes through a beam-splitter $S_{\mathrm{BS}}=\left(\begin{smallmatrix}\cos(\theta)&\sin(\theta)\\ \sin(\theta)&-\cos(\theta)\end{smallmatrix}\right)$, which corresponds to the coordinate transformations $\left(x_{A}^{\prime},x_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(x_{A},x_{B}\right)^{T}$, $\left(p_{A}^{\prime},p_{B}^{\prime}\right)^{T}=S_{\mathrm{BS}}\left(p_{A},-p_{B}\right)^{T}$, we obtain the Wigner function of one output mode of the beam-splitter,
$W_{B}^{\prime}(x_{B}^{\prime},p_{B}^{\prime})$ $=\frac{1}{2\pi\left(-2c\sin(\theta)\cos(\theta)+m\cos^{2}(\theta)+n\sin^{2}(\theta)\right)}$
$\times\exp\left(-\frac{x_{B}^{\prime 2}+p_{B}^{\prime 2}}{2\left(-2c\sin(\theta)\cos(\theta)+m\cos^{2}(\theta)+n\sin^{2}(\theta)\right)}\right).$ (35)
@@
IV. Detailed proof of criterion (I)
@@
Let $0\leqslant\theta<\pi$ be a constant angle ($\theta\neq\pi/2$), and let $(X,P)$ represent a constant coordinate. Then, define the following two coordinate transformations according to $(x^{\prime},p^{\prime})$, along with the corresponding state transformations $h_{1},h_{2}$,
@@
where in the last row we have applied the substitution $\left(x\rightarrow x+\frac{\sin\theta}{\cos\theta}X,p\rightarrow p+\frac{\sin\theta}{\cos\theta}P\right)$. Now, assume that $\rho_{AB}$ is a separable pure state, meaning $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$ should also correspond to a physical state. Next, perform a beam-splitter transformation given by $\left(\begin{smallmatrix}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{smallmatrix}\right)$ on the two modes of $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$. The resulting Wigner function is given by
@@
V. Criterion (I) is sufficient and necessary for Gaussian states
@@
$W_{AB}(\cos\theta x,\cos\theta p,\sin\theta x^{\prime},\sin\theta p^{\prime})=\frac{1}{4\pi^{2}\sqrt{\left(c_{1}^{2}-mn\right)\left(c_{2}^{2}-mn\right)}}\times$ (43)
[…]
$\left.\left.-c_{1}t\sin(2\theta)\left(c_{1}c_{2}p^{2}+mnt^{2}x^{2}\right)\right]/\left(2t^{2}\left(c_{1}^{2}-mn\right)\left(mn-c_{2}^{2}\right)\right)\right\}$
@@
$I$ $=\iint_{-\infty}^{\infty}W_{AB}(\cos\theta x,\cos\theta p,\sin\theta x^{\prime},\sin\theta p^{\prime})\text{d}x\text{d}p$ (44)
$=\frac{1}{2\pi\sqrt{\frac{\left(m+nt^{2}+\cos(2\theta)(m-nt^{2})-2c_{1}t\sin(2\theta)\right)\left(mt^{2}+n+\cos(2\theta)(mt^{2}-n)+2c_{2}t\sin(2\theta)\right)}{4t^{2}}}}.$
@@
$t^{2}$ $=\frac{\left|c_{2}\right|m+\left|c_{1}\right|n}{\left|c_{1}\right|m+\left|c_{2}\right|n},$ (46)
$\tan(2\theta)$ $=\frac{-2\sqrt{\left|c_{2}\right|m+\left|c_{1}\right|n}\sqrt{\left|c_{1}\right|m+\left|c_{2}\right|n}}{(m-n)(m+n)}.$
By simplifying the entanglement condition $I>1/(2\pi)$, we arrive precisely at Eq. (42). Note that when $m=n$, the optimum is reached in the limit $\theta\to\pi/4$, and when $c_{1}=c_{2}=0$ the state is separable and $I\leq 1/(2\pi)$. Both cases remain consistent with Eq. (42).
VI. A comparison between criterion (III) and previous criteria based on Wigner functions
Reference [11] introduces an entanglement criterion based on reduced states. The Wigner function is first expressed in the $\left\{\hat{a}_{+},\hat{a}_{-}\right\}$ basis as $W_{\rho}\left(\alpha_{+},\alpha_{-}\right)$. By tracing out the $a_{-}$ mode, the reduced Wigner function is obtained as $W_{\text{tr}_{-}(\rho)}\left(\alpha_{+}\right)=\int\text{d}^{2}\alpha_{-}W_{\rho}\left(\alpha_{+},\alpha_{-}\right)$. If any point in the reduced Wigner function $W_{\text{tr}_{-}(\rho)}$ is negative, the state must be entangled. This criterion is a special case of our lower bound when choosing $(x^{\prime}=x_{0}-x,p^{\prime}=p_{0}-p)$, where $x_{0}=\sqrt{2}x_{+}$ and $p_{0}=\sqrt{2}p_{+}$ are constants. This is because